# A Systematic Survey on Large Language Models for Algorithm Design

### v1

@@ 6.1.2. Continuous Optimization
Here we delve into works that deal with optimization problems featuring continuous design variables, encompassing both single- and multi-objective optimization. Specific tasks or benchmarks include synthetic functions as well as real-world applications.
In the realm of single-objective optimization, LLaMEA (van Stein and Bäck, 2024) utilizes LLMs to automate the evolution of algorithm design. It demonstrates the effectiveness in generating new metaphor-based optimization algorithms on BBOB benchmark (Hansen and Ros, 2010) within IOHexperimenter benchmarking tool (de Nobel et al., 2024), which supports evaluating the quality of the generated algorithms and also provides feedback to the LLM during evolution. Instead of creating new algorithms, EvolLLM (Lange et al., 2024) introduces a prompt strategy that enables LLM-based optimization to act as an Evolution Strategy (ES) and showcases robust performance on synthetic BBOB functions and neuroevolution tasks. OPRO (Yang et al., 2024) illustrates that LLMs effectively capture optimization directions for linear regression problems by leveraging the past optimization trajectory from the meta-prompt. Additionally, LEO (Brahmachary et al., 2024) devises an explore-exploit policy using LLMs for solution generation, the method has been tested and validated in both single- and multi-objective benchmark functions, as well as industrial engineering problems including supersonic nozzle shape optimization, heat transfer, and windfarm layout optimization. In a slight departure from the aforementioned methods that directly employ LLMs for solution generation, LAEA (Hao et al., 2024) introduces LLM-based surrogate models for regression and classification tasks to assess the quality of new solutions for selection. This approach has been validated on 2D test functions using nine mainstream LLMs.
Compared to single-objective continuous optimization, multi-objective optimization problems (MOPs) seek to identify a set of optimal solutions, referred to as Pareto optimal solution sets. Multi-Objective Evolutionary Algorithms (MOEAs) serve as effective tools for addressing MOPs across various domains. An initial exploration of utilizing LLMs to tackle MOPs is introduced in MOEA/D-LMO (Liu et al., 2023a). Benefiting from the decomposition-based framework, the in-context learning process of LLMs is easily incorporated to generate candidate solutions for each subproblem derived from the original MOP. In the realm of large-scale MOPs, LLM-MOEA* (Singh and Bali, 2024) showcases the inferential capabilities of LLMs in multi-objective sustainable infrastructure planning problem. The study highlights the LLM’s proficiency in filtering crucial decision variables, automatically analyzing the Pareto front, and providing customized inferences based on varying levels of expertise. Additionally, CMOEA-LLM (Wang et al., 2024b) leverages LLMs with evolutionary search operators to address the challenges of constrained MOPs and exhibits robust competitiveness in DAS-CMOP test suite (Fan et al., 2020).
@@ 6.1.3. Bayesian Optimization
Bayesian optimization (BO) is a model-based optimization paradigm for solving expensive optimization problems and has found wide application in various real-world scenarios (Frazier and Wang, 2016). It typically employs a surrogate model to approximate the expensive function and well-designed Acquisition Functions (AFs) to select potential solutions in a sample-efficient manner. To facilitate the direct generation of solutions using LLMs, HPO-LLM* (Zhang et al., 2023) provides LLMs with an initial set of instructions that outlines the specific dataset, model, and hyperparameters to propose recommended hyperparameters for evaluation in Hyperparameter Optimization (HPO) tasks. This approach has demonstrated comparably or better performance when compared with random search and Bayesian optimization on HPOBench dataset (Eggensperger et al., 2021). Furthermore, LLAMBO incorporates LLM capabilities to enhance the efficiency of model-based BO, in which three specific enhancements throughout the BO pipeline including zero-shot warmstarting, surrogate models and candidate point sampler, as well as the end-to-end performance have been systematically investigated, the evaluated tasks are selected from Bayesmark (Uber, 2020) and HPOBench (Eggensperger et al., 2021). Instead of utilizing LLM for direct solution generation, BO-LIFT* (Ramos et al., 2023) utilizes predictions with uncertainties provided by a Language-Interfaced Fine-Tuning (LIFT) framework (Dinh et al., 2022) with LLMs to perform BO for catalyst optimization using natural language. EvolCAF (Yao et al., 2024a) introduces a novel paradigm to design acquisition functions automatically for cost-aware Bayesian optimization, the approach showcases remarkable efficiency and generalization on synthetic functions and hyperparameter tuning task in JAHS-Bench-201 data set (Bansal et al., 2022). FunBO (Aglietti et al., 2024) discoveres novel and well-performing AFs for BO by extending FunSearch (Romera-Paredes et al., 2024), the discovered AFs are evaluated on various global optimization benchmarks in and out of the training distribution and HPO tasks for RBF-based SVM and AdaBoost algorithms.
@@ 6.1.4. Prompt Optimization
Prompt optimization aims to identify the most effective task prompt that maximizes the performance of the LLM on a specific task dataset. Previous studies have primarily focused on discrete optimization for natural language prompts (Deng et al., 2022; Zhang et al., 2022; Xu et al., 2022; Prasad et al., 2022) or continuous optimization in the embedding space (Li and Liang, 2021; Lester et al., 2021). Nevertheless, despite of requiring specialized training for each specific task, these approaches typically necessitate access to the logits or internal states of LLMs, which may not be applicable when the LLM can only be accessed through an API. To address these issues, recent works propose to model the optimization problem in natural language with LLMs as prompts. APE (Zhou et al., 2022) utilizes LLM as an inference model to generate instruction candidates directly based on a small set of demonstrations in the form of input-output pairs. This approach has demonstrated human-level performance on various tasks, including Instruction Induction (Honovich et al., 2022) and Big-Bench Hard (BBH) (Suzgun et al., 2022). OPRO (Yang et al., 2024) enables the LLM as an optimizer to gradually generate new prompts based on the full optimization trajectory, the optimizer prompt showcases significant improvement compared with human-designed prompts on GSM8K (Cobbe et al., 2021) and BBH. Inspired by the numerical gradient descent method, APO (Pryzant et al., 2023) conducts textual “gradient descent” by identifying the current prompts’ flaws and adjusting the prompt in the opposite semantic direction of the gradient. The outcomes showcase substantial enhancements with no hyperparameter tuning or model training on four NLP benchmark classification tasks, namely Jailbreak, Ethos (Mollas et al., 2020), Liar (Wang, 2017) and Sarcasm (Farha and Magdy, 2020). Similarly, Tang et al. (2024c) introduce a gradient-inspired LLM-based prompt optimizer named GPO and evaluate its performance across complex reasoning, knowledge-intensive, and common NLP tasks on datasets including BBH, GSM8K, MMLU (Hendrycks et al., 2020), WSC (Levesque et al., 2012) and WebNLG (Gardent et al., 2017). Guo et al. (2024b) present a collaborative optimization approach that integrates a gradient-based optimizer and an LLM-based optimizer in an interleaved fashion, the combined method achieves consistent improvements compared with baseline prompt tuning methods on NLU and image classification tasks. By harnessing the benefits of evolutionary algorithms that that strike a good balance between exploration and exploitation, Guo et al. (2023) introduce a discrete prompt tuning framework named EvoPrompt that prompts LLM to act like evolutionary operators in generating new candidate prompts. The superiority of EvoPrompt over manually designed prompts and previous automatic prompt generation approaches has been demonstrated across 31 datasets, encompassing tasks related to language understanding, generation, and the BIG-Bench benchmark.
@@ 6.1.5. Optimization Modeling
Optimization modeling (Berry and Houston, 1995) aims to construct a mathematical model of a real-world optimization problem in a standard form represented with an objective function and a set of constraints (AhmadiTeshnizi et al., 2023; Tang et al., 2024a). In addition to designing effective algorithms to solve optimization problems, the process of modeling or formulating the problems also presents a significant challenge. Even when employing the same optimization algorithm, different modeling formulation may lead to significant differences in solving efficiency. Besides, creating an efficient formulation often demands expertise in optimization skills and specialized domain knowledge in a wide range of applications (Wasserkrug et al., 2024; Saghafian et al., 2015; Shakoor et al., 2016).
Recently, LLMs have opened up promising avenues for automating optimization modeling, facilitating the formulation of problems and the implementation of solutions, thereby democratizing expert knowledge. AhmadiTeshnizi et al. (2023) propose OptiMUS, an LLM-based agent designed to formulate and solve optimization problems interpreted from natural language. The performance of the method is evaluated on a proposed novel dataset named NLP4LP (AhmadiTeshnizi et al., 2023) (Natural Language Processing for Linear Programming), which includes 41 linear programming (LP) and 11 mixed integer linear programming (MILP) problems. Despite its outstanding performance, the heavy reliance on sensitive data submission to proprietary LLMs can pose data privacy concerns in industrial applications. To solve this issue, Tang et al. (2024a) propose training open-source LLMs for optimization modeling that are applicable to real-world industries. They develop and implement OR-Instruct, a semi-automated process for generating synthetic data customized to meet specific requirements. The best-performing resulting model, ORLM, demonstrates state-of-the-art performance on NL4Opt (Ramamonjison et al., 2023), MAMO (Huang et al., 2024a), and the proposed first industrial benchmark, IndustryOR. Ahmed and Choudhury (2024) develop a progressive fine-tuning framework named LM4OPT to enhance LLMs’ specificity for optimization problem formulation task on NL4Opt dataset. Wasserkrug et al. (2024) develop Decision Optimization CoPilot (DOCP), which is accessible to a broad range of decision-makers lacking problem-dependent modeling skills. This tool aids in decision-making by interacting with the user in natural language and utilizes human’s knowledge and feedback to create and solve a problem-specific optimization model. The experiments are conducted targeting three presented requirements for DOCP on facility location problems.
@@ 6.1.6. Other Applications
In addition to optimization applications that can be categorized into various fields as mentioned above, we have also discovered applications of LLMs in other optimization-related domains, which may intersect to varying degrees with the aforementioned classifications. For ease of reference, we have categorized them as other applications.
@@
LAEA (Hao et al., 2024) FS, CoT LLMaP Ellipsoid, Rosenbrock, Ackley, Griewank
@@
GPO (Tang et al., 2024c) LLMaO FS BBH, GSM8K, MMLU, WSC, WebNLG
MaaO (Guo et al., 2024b) LLMaO FS NLU Tasks, Image Classification Tasks
@@ 6.2. Machine Learning
L-AutoDA (Guo et al., 2024c) LLMaD FS, OS, CoT Adversarial Attack
In this subsection, we investigate the applications of LLMs in the machine learning domain, focusing on their contribution to algorithmic design. Our analysis begins with a focus on significant application areas, notably task planning and reinforcement learning—key benchmarks in the pursuit of generalized artificial intelligence. Subsequently, we probe into specialized applications including neural architecture search, graph-based learning, and the automation of dataset labeling. While it is not feasible to discuss every application exhaustively, we feature additional uses of LLMs in tasks like tensor network structure search and adversarial attack algorithm generation. These applications are summarized in Table 3.
@@ 6.2.1. Task Planning with LLMs
The advent of LLMs has garnered significant interest in their application to planning tasks, owing to their remarkable generative capabilities of instructions in the form of natural language. Pioneering work by Huang et al. (2022) demonstrates the effectiveness of LLMs in leveraging world knowledge to produce actionable plans for VirtualHome environments, showcasing their potentiality in such contexts. This method was further enhanced by Ahn et al. (2022), who employed a value function to ground the world knowledge, enabling the derivation of an optimal policy for robotic task execution in lifelike kitchen settings. In subsequent research, Huang et al. (2023) broaden the scope of grounded information in a more adaptive way. Moreover, Lin et al. (2023) utilize greedy search in combination with heuristic algorithms to refine the planning system. Distinctly, Singh et al. (2023) tackle planning challenges by integrating programming languages, thereby generating code to streamline planning processes. A more comprehensive approach, named SayCanPay (Hazra et al., 2024), is developed for planning by considering both the affordance function and reward within the task framework.
6.2.2. Reinforcement Learning.
@@ 6.2.3. Neural Architecture Search
Neural Architecture Search (NAS) has emerged as a significant focus within the AutoML community, with recent enhancements attributed to the capabilities of LLMs. LLMs are most commonly used for designing neural architectures. For example, Chen et al. (2024a) have integrated LLMs with evolutionary search to successfully generate code for diverse tasks, including NAS for the MNIST dataset. Nasir et al. (2023) introduce a quality-diversity algorithm tailored to NAS, producing architectures for CIFAR-10 and NAS-bench-201 benchmarks. Moreover, Morris et al. (2024) introduce guided evolution for the development of neural architectures and has been the first to suggest the concept of the evolution of thought in algorithm design. LLMs have been also used for predicting the performance of the network. Jawahar et al. (2023) employ LLMs in predicting NAS performance, combining this approach with evolutionary search to effectively create novel network architectures. Different from the LLM-based architecture design and score prediction, (Zhou et al., 2024a) investigate how to adopt LLM for design principle transfer to reduce and direct the search space. A notable search efficiency improvement is achieved when compared to existing transfer methods for NAS.
@@ 6.2.4. Graph Learning
Graph learning constitutes a significant field within machine learning. With the advancing capabilities of LLMs in symbolic reasoning and graph processing, their integration into graph learning has emerged as a recent trend. For example, Chen et al. (2023b) apply LLMs to the task of labeling in Text-Attributed Graphs (TAGs), capitalizing on the language task proficiency of LLMs. Both Mao et al. (2024) and Chen et al. (2024c) adopt LLMs in an evolutionary framework for designing functions. The former evolves heuristic code functions to identify critical nodes in a graph while the latter identifies meta-structures within heterogeneous information networks to enhance their interpretability. Moreover, knowledge graphs have also seen substantial benefits from the application of LLMs. Zhang et al. (2024d) introduce AutoAlign, a method that employs LLMs to semantically align entities across different knowledge graphs, and Feng et al. (2023b) develop the knowledge search language, which leverages LLMs to effectively conduct searches within knowledge graphs.
@@ 6.2.5. Dataset Labeling
LLMs have been used for mining semantic and multi-modal information from datasets. LLMs are emplyed to train interpretable classifiers to extract attributes from images (Chiquier et al., 2024) and to generate label functions for weakly supervised learning (Guan et al., 2023).

### v2

@@ 6.1.2. Continuous Optimization
In the realm of continuous single-objective optimization, LLaMEA (van Stein and Bäck, 2024) utilizes LLMs to automate the evolution of algorithm design. It demonstrates the effectiveness in generating new metaphor-based optimization algorithms on BBOB benchmark (Hansen and Ros, 2010) within IOHexperimenter benchmarking tool (de Nobel et al., 2024), which supports evaluating the quality of the generated algorithms and also provides feedback to the LLM during evolution. Instead of creating new algorithms, EvolLLM (Lange et al., 2024) introduces a prompt strategy that enables LLM-based optimization to act as an Evolution Strategy (ES) and showcases robust performance on synthetic BBOB functions and neuroevolution tasks. OPRO (Yang et al., 2024) illustrates that LLMs can effectively capture optimization directions for linear regression problems by leveraging the past optimization trajectory from the meta-prompt. Additionally, LEO (Brahmachary et al., 2024) devises an explore-exploit policy using LLMs for solution generation, the method has been tested and validated in both benchmark functions as well as industrial engineering problems. Different with directly employing LLMs for generating solutions, LAEA (Hao et al., 2024) introduces LLM-based surrogate models for regression and classification tasks to assess the quality of new solutions for selection. This approach has been validated on 2D test functions using nine mainstream LLMs.
Multi-objective optimization problems (MOPs) seek to identify a set of optimal solutions, referred to as Pareto optimal solution set. An initial exploration of utilizing LLMs to tackle MOPs is introduced in MOEA/D-LMO (Liu et al., 2023a). Benefiting from the decomposition-based framework, the in-context learning process of LLMs is easily incorporated to generate candidate solutions for each subproblem derived from the original MOP. In the realm of large-scale MOPs, LLM-MOEA* (Singh and Bali, 2024) showcases the inferential capabilities of LLMs in multi-objective sustainable infrastructure planning problem. The study highlights the LLM’s proficiency in filtering crucial decision variables, automatically analyzing the Pareto front, and providing customized inferences based on varying levels of expertise. Additionally, CMOEA-LLM (Wang et al., 2024b) leverages LLMs with evolutionary search operators to address the challenges of constrained MOPs and exhibits robust competitiveness in DAS-CMOP test suite (Fan et al., 2020).
@@ 6.1.3. Bayesian Optimization
Bayesian optimization (BO) is a model-based optimization paradigm for solving expensive optimization problems and has found wide application in various real-world scenarios (Frazier and Wang, 2016). It typically employs a surrogate model to approximate the expensive function and well-designed Acquisition Functions (AFs) to select potential solutions in a sample-efficient manner. To facilitate the direct generation of solutions using LLMs, HPO-LLM* (Zhang et al., 2023) provides LLMs with an initial set of instructions that outlines the specific dataset, model, and hyperparameters to propose recommended hyperparameters for evaluation in Hyperparameter Optimization (HPO) tasks. Furthermore, LLAMBO (Liu et al., 2024a) incorporates LLM capabilities to enhance the efficiency, in which three specific enhancements throughout the BO pipeline have been systematically investigated on tasks selected from Bayesmark (Uber, 2020) and HPOBench (Eggensperger et al., 2021). Instead of utilizing LLMs for direct solution generation, BO-LIFT* (Ramos et al., 2023) utilizes predictions with uncertainties provided by a Language-Interfaced Fine-Tuning (LIFT) framework (Dinh et al., 2022) with LLMs to perform BO for catalyst optimization using natural language. EvolCAF (Yao et al., 2024a) introduces a novel paradigm to design AFs automatically for cost-aware BO, the approach showcases remarkable efficiency and generalization for synthetic functions and HPO tasks on JAHS-Bench-201 (Bansal et al., 2022). FunBO (Aglietti et al., 2024) discoveres novel and well-performing AFs for BO by extending FunSearch (Romera-Paredes et al., 2024), the discovered AFs are evaluated on various global optimization benchmarks in and out of the training distribution and HPO tasks for RBF-based SVM and AdaBoost algorithms.
@@ 6.1.4. Prompt Optimization
Prompt optimization aims to identify the most effective task prompt that maximizes the performance of the LLM on a specific task dataset. Despite of requiring specialized training for each specific task, traditional discrete or continuous approaches (Prasad et al., 2022; Li and Liang, 2021) typically necessitate access to the logits or internal states of LLMs, which may not be applicable when the LLM can only be accessed through an API. To address these issues, recent works propose to model the optimization problem in natural language with LLMs as prompts. APE (Zhou et al., 2022) utilizes the LLM as an inference model to generate instruction candidates directly based on a small set of demonstrations in the form of input-output pairs. This approach has demonstrated human-level performance on various tasks, including Instruction Induction (Honovich et al., 2022) and Big-Bench Hard (BBH) (Suzgun et al., 2022). OPRO (Yang et al., 2024) enables the LLM as an optimizer to gradually generate new prompts based on the full optimization trajectory, the optimizer prompt showcases significant improvement compared with human-designed prompts on GSM8K (Cobbe et al., 2021) and BBH. Inspired by the numerical gradient descent method, APO (Pryzant et al., 2023) conducts textual “gradient descent” by identifying the current prompts’ flaws and adjusting the prompt in the opposite semantic direction of the gradient. Similar practices are also found in the gradient-inspired LLM-based prompt optimizer named GPO (Tang et al., 2024c), as well as the collaborative optimization approach (Guo et al., 2024a) integrating a gradient-based optimizer and an LLM-based optimizer. Differently, Guo et al. (2023) introduce a discrete prompt tuning framework named EvoPrompt that prompts LLM to act like evolutionary operators in generating new candidate prompts, harnessing the benefits of evolutionary algorithms that strike a good balance between exploration and exploitation.
@@ 6.1.5. Optimization Modeling
Optimization modeling (Berry and Houston, 1995) aims to construct a mathematical model of a real-world optimization problem in a standard form, represented with an objective function and a set of constraints (AhmadiTeshnizi et al., 2023; Tang et al., 2024a). In addition to resulting significant differences in solving efficiency, creating an efficient formulation often demands expertise in optimization skills and specialized domain knowledge in various applications. Recently, LLMs have opened up promising avenues for automating optimization modeling, facilitating the formulation of problems and the implementation of solutions, thereby democratizing expert knowledge. AhmadiTeshnizi et al. (2023) propose OptiMUS, an LLM-based agent designed to formulate and solve optimization problems interpreted from natural language. The performance of the method is evaluated on a proposed novel dataset named NLP4LP (AhmadiTeshnizi et al., 2023), which includes 41 linear programming and 11 mixed integer linear programming problems. Despite its outstanding performance, the heavy reliance on sensitive data submission to proprietary LLMs can pose data privacy concerns in industrial applications.
To solve this issue, Tang et al. (2024a) propose training open-source LLMs for optimization modeling that are applicable to real-world industries. They develop and implement OR-Instruct, a semi-automated process for generating synthetic data customized to meet specific requirements. The best-performing resulting model, ORLM, demonstrates state-of-the-art performance on NL4Opt (Ramamonjison et al., 2023), MAMO (Huang et al., 2024a), and the proposed IndustryOR benchmark.
Ahmed and Choudhury (2024) develop a progressive fine-tuning framework named LM4OPT to enhance LLMs’ specificity for optimization problem formulation task on NL4Opt dataset. Wasserkrug et al. (2024) develop Decision Optimization CoPilot (DOCP), which is accessible to a broad range of decision-makers lacking problem-dependent modeling skills. This tool aids in decision-making by interacting with the user in natural language and utilizes human’s knowledge and feedback to create and solve a problem-specific optimization model. The experiments are conducted targeting three presented requirements for DOCP on facility location problems.
@@ 6.1.6. Other Applications
In addition to optimization applications that can be categorized into various domains as mentioned above, we have also discovered applications of LLMs in other optimization-related fields, which may intersect to varying degrees with the aforementioned classifications. For ease of reference, we have categorized them as ”other applications” and will introduce them separately below.
@@
LAEA (Hao et al., 2024) LLMaP FS, CoT Ellipsoid, Rosenbrock, Ackley, Griewank
@@
GPO (Tang et al., 2024c) LLMaO FS Reasoning, Knowledge-intensive, NLP Tasks
MaaO (Guo et al., 2024a) LLMaO FS NLU Tasks, Image Classification Tasks
@@ 6.2. Machine Learning
L-AutoDA (Guo et al., 2024b) LLMaD FS, OS, CoT Adversarial Attack
In this subsection, we investigate the applications of LLMs in the machine learning domain, focusing on their contribution to algorithmic design. These applications are summarized in Table 2.
@@ 6.2.1. Task Planning with LLMs
The advent of LLMs has garnered significant interest in their application to planning tasks, owing to their remarkable generative capabilities of instructions in the form of natural language. Pioneering work by Huang et al. (2022) demonstrates the effectiveness of LLMs in leveraging world knowledge to produce actionable plans for VirtualHome environments. This method was further enhanced by Ahn et al. (2022), who employed a value function to ground the world knowledge, enabling the derivation of an optimal policy for robotic task execution in lifelike kitchen settings. In subsequent research, Huang et al. (2023) broaden the scope of grounded information in a more adaptive way. Moreover, Lin et al. (2023) utilize greedy search in combination with heuristic algorithms to refine the planning system. Distinctly, Singh et al. (2023) tackle planning challenges by integrating programming languages, thereby generating code to streamline planning processes. A more comprehensive approach, named SayCanPay (Hazra et al., 2024), is developed for planning by considering both the affordance function and reward within the task framework.
6.2.2. Reinforcement Learning
@@ 6.2.3. Neural Architecture Search
Neural Architecture Search (NAS), which is a significant focus within the AutoML community, has been investigated in many LLM4AD papers. For example, Chen et al. (2024a) have integrated LLMs with evolutionary search to successfully generate NAS code for diverse tasks. Nasir et al. (2023) introduce a quality-diversity algorithm tailored to NAS, producing architectures for CIFAR-10 and NAS-bench-201 benchmarks. Moreover, Morris et al. (2024) introduce guided evolution for the development of neural architectures and suggest the concept of the evolution of thought in algorithm design. Except for using LLM for design, Jawahar et al. (2023) employ LLMs in predicting NAS performance, combining this approach with evolutionary search to effectively create novel network architectures. In contrast to the LLM-based architecture design and performance prediction, Zhou et al. (2024a) explore the adoption of LLMs for transferring design principles to narrow and guide the search space.
@@ 6.2.4. Graph Learning
Graph learning is another application with the advancing capabilities of LLMs in symbolic reasoning and graph processing. For example, Chen et al. (2023b) apply LLMs to the task of labeling in Text-Attributed Graphs (TAGs), capitalizing on the language task proficiency of LLMs. Both Mao et al. (2024) and Chen et al. (2024c) adopt LLMs in an evolutionary framework for designing functions. The former evolves heuristic code functions to identify critical nodes in a graph while the latter identifies meta-structures within heterogeneous information networks to enhance their interpretability. Moreover, knowledge graphs have also seen substantial benefits from the application of LLMs. Zhang et al. (2024d) introduce AutoAlign, a method that employs LLMs to semantically align entities across different knowledge graphs, and Feng et al. (2023b) develop the knowledge search language to effectively conduct searches within knowledge graphs.
@@ 6.2.5. Dataset Labeling
LLMs have been used for mining semantic and multi-modal information from datasets. LLMs are employed to train interpretable classifiers to extract attributes from images (Chiquier et al., 2024) and to generate label functions for weakly supervised learning (Guan et al., 2023).